# Mining Limited Data Sufficiently: A BERT-inspired Approach for CSI Time Series Application in Wireless Communication and Sensing

### v1

@@ Abstract
Channel State Information (CSI) is the cornerstone in both wireless communication and sensing systems. In wireless communication systems, CSI provides essential insights into channel conditions, enabling system optimizations like channel compensation and dynamic resource allocation. However, the high computational complexity of CSI estimation algorithms necessitates the development of fast deep learning methods for CSI prediction. In wireless sensing systems, CSI can be leveraged to infer environmental changes, facilitating various functions, including gesture recognition and people identification. Deep learning methods have demonstrated significant advantages over model-based approaches in these fine-grained CSI classification tasks, particularly when classes vary across different scenarios. However, a major challenge in training deep learning networks for wireless systems is the limited availability of data, further complicated by the diverse formats of many public datasets, which hinder integration. Additionally, collecting CSI data can be resource-intensive, requiring considerable time and manpower.
To address these challenges, we propose CSI-BERT2 for CSI prediction and classification tasks, effectively utilizing limited data through a pre-training and fine-tuning approach. Building on CSI-BERT1, we enhance the model architecture by introducing an Adaptive Re-Weighting Layer (ARL) and a Multi-Layer Perceptron (MLP) to better capture sub-carrier and timestamp information, effectively addressing the permutation-invariance problem. Furthermore, we propose a Mask Prediction Model (MPM) fine-tuning method to improve the model’s adaptability for CSI prediction tasks. Experimental results demonstrate that CSI-BERT2 achieves state-of-the-art performance across all tasks with relatively fast computation speeds. To facilitate future research, we will make our code and dataset publicly available upon publication.
@@ I Introduction
Channel State Information (CSI) plays a critical role in both wireless sensing and wireless communication systems by capturing the propagation characteristics of the wireless channel. This information provides valuable insights into the radio environment, including the channel conditions, signal quality, and the motion and location of objects. In wireless communication systems, CSI is essential for optimizing various processes such as channel compensation, adaptive modulation and coding, user selection and scheduling. By effectively utilizing CSI, these systems can achieve better performance, increased spectrum efficiency, and more reliable communication, ultimately enhancing the overall user experience [1, 2, 3]. In wireless sensing applications, CSI can enhance the ability to detect and track objects, leading to improved environmental awareness. With the advantage of privacy protection, low cost, and penetration ability, wireless sensing has been widely used in many areas like fall detection [4], localization [5], and people identification [6].
Despite the importance of CSI in wireless systems, it is challenging to acquire, process, and utilize CSI data. For example, many deep learning methods in wireless systems require substantial data for training, but the data available in wireless communication is often limited, and public datasets may have different formats, making it challenging to use them together [6]. Additionally, collecting CSI datasets on one’s own can be a difficult task.
A common solution for data scarcity is pre-training, which has been widely utilized in fields such as Natural Language Processing (NLP) [7, 8], Computer Vision (CV) [9, 10], and Music Information Retrieval (MIR) [11, 12]. The pre-training process typically employs a self-supervised approach, enabling models to learn the underlying structure and patterns of the data. This can enhance the model’s performance on downstream tasks, as well as improve its convergence speed and generalization capacity, particularly in scenarios with limited training data [13, 14]. Pre-training not only allows models to learn and understand the data more effectively, but it also enables the use of additional unlabeled data for training. Currently, pre-training is not exclusively applied to address data scarcity. In fields such as NLP and CV, as datasets continue to grow, pre-training has emerged as a powerful strategy to bolster model generalization capacity [15, 16]. Moreover, even when the inference scenario differs from the training scenario, models can quickly adapt to new contexts through transfer learning methods [17, 18], such as fine-tuning.
Even though, pre-training and fine-tuning models on different datasets is uncommon in certain specialized fields, such as MIR and signal processing, which often face challenges due to limited data. On one hand, some tasks lack additional data for pre-training. On the other hand, simply fine-tuning a pre-trained model on a new dataset may result in poor performance when the pre-training data is insufficient. Consequently, most research in these fields opts to pre-train and fine-tune models using the same dataset [19]. The unsupervised learning during the pre-training phase effectively helps models extract valuable insights from the limited data available.
Bidirectional Encoder Representations from Transformers (BERT) [20] is a pioneering work that first explored the efficiency of pre-training method in NLP. With the help of its bi-directional attention mechanism, which efficiently extracts the relationship of context, and the Mask Language Model (MLM) pre-training task, which helps the model learn the basic data structure, BERT has become the foundation for current understanding models. Since wireless information sequences also exhibit contextual relationships like natural language, BERT can also be used to extract their inherent features. Previous researchers have explored the application of BERT in wireless sensing tasks, including localization [21, 22] and radio maps construction [23]. However, most of these studies directly convert continuous wireless signals into discrete tokens, resulting in significant information loss. Additionally, their approach simply adapted BERT from NLP without any costomized design, which is not suitable due to the different structure of wireless signals and natural language.
In contrast, we have previously proposed a model named CSI-BERT1 [24] for CSI recovery. By redesigning the structure of BERT, we make it more suitable for CSI sequences, such as using a linear layer-based token embedding to replace the traditional tokenization method and designing a time embedding for CSI time series. While CSI-BERT1 has shown promising performance, we have identified some shortcomings, such as the permutation invariance of the time embedding [25] and the gap between training and inference.
In this paper, we address these issues by optimizing both the model structure and the training process, as illustrated in Fig. 1. We further introduce CSI-BERT2 for CSI prediction and classification. By incorporating a novel embedding layer that effectively captures the relationships between subcarriers and timestamps, our model demonstrates a robust capability for processing CSI time series. Additionally, by designing and utilizing the [MASK] token within the CSI series, we enable our model to adapt to various wireless scenarios, as shown in Fig. 2. We evaluate our model on two publicly available datasets, WiGesture [24] and WiFall [26], along with a self-collected dataset, WiCount. The experimental results indicate that CSI-BERT2 achieves state-of-the-art performance across all tasks. In summary, the main contributions of this work are:
@@
(1) A Multifunctional Framework for CSI Time Series: We propose CSI-BERT2, a multifunctional model for CSI time series applications in wireless systems, including CSI prediction and CSI classification. To realize these, we optimize the structure and training process of the original CSI-BERT1 model.
@@
(2) Improved Model Structure: We design a novel time embedding technique that using Multi-Layer Perceptron (MLP) to further encode time information, to solve the permutation invariant problem of the encoder structure, which can be detrimental for time series. Second, we use a simple attention mechanism called Adaptive Re-Weighting Layer (ARL) to better capture the relationships between subcarriers.
(3) Improved Training Process: To make full use of the limited data, we first pre-train the model in a unsupervised manner by MLM task and then fine-tune it in specific downstreaming tasks. Specifically, we propose a Mask Prediction Model (MPM) fine-tune method to make the encoder structure more suitable for prediction tasks.
(4) Experiment Evaluation: Experimental results demonstrate that our CSI-BERT2 model outperforms all other models across all tasks. In the CSI prediction task, it achieves the lowest error while maintaining relatively fast computation speeds. In the CSI classification task, it attains the highest accuracy in various applications, including gesture recognition, people identification, fall detection, and crowd counting. Notably, it can directly utilize discontinuous data caused by packet loss and effectively handle scenarios where the testing set has a different sampling rate than the training set, even when the training set itself has varying sampling rates.
@@
The structure of this paper is organized as follows. In Section II, we first introduce the fundamental principles of CSI, along with the background on CSI prediction and classification. We also review relevant works in these areas. In Section III, we provide a detailed description of the model structure and the training process for different tasks. In Section IV, we present a series of comparative experiments to demonstrate the advantages of our method, along with several ablation studies to illustrate the efficiency of our design. Finally, in Section V, we summarize the paper and highlight potential directions for future research.
(a) CSI-BERT1 Framework [24]
(b) CSI-BERT2 Framework
Figure 1: Architecture of Two Generations of CSI-BERT: In Fig. 1b, ‘T’ represents token, ’t’ represents time stamp, and ’P’ represents the position.
(a) CSI Prediction
(b) CSI Classification
Figure 2: Two Tasks in This Paper: (a) The CSI prediction task focuses on swiftly predicting future CSI series for proactive communication optimization. (b) The CSI classification task utilizes CSI for specific sensing applications, such as gesture recognition. Notably, our CSI-BERT2 can efficiently process CSI data at varying sampling rates, a challenge for most other models.
II Preliminary
@@ II-A CSI Basics
Following [27], in Wi-Fi communication, CSI characterizes the way Wi-Fi signals propagate from the transmitter (TX) to the receiver (RX). The amplitude and phase of CSI are influenced by multipath effects, which include amplitude attenuation and phase shift. Each element of the CSI Matrix $H$ can be given by:
@@ II-B CSI Prediction
In wireless communication, obtaining the CSI matrix is a challenging task, especially in future mmWave scenarios [28]. Considering the large number of antennas, the dimension of the CSI matrix will also be substantial, significantly increasing the complexity of the channel estimation algorithms. On the other hand, when objects in the environment are moving at high speeds, the channel coherence time will be greatly reduced, thereby increasing the requirements for channel estimation time. Nevertheless, shown as Fig. 2a, CSI is crucial for optimizing wireless communication systems, making rapid estimation essential. Moreover, if we can predict future CSI in advance, we will have more time available for system optimizations, such as channel compensation.
Recently, some works have explored the use of additional information to predict CSI. In [29], the authors used multimodal sensory data, including received pilots, user’s position, and previous downlink CSI to predict the current CSI. In [30], a deep learning-based framework using a 3-D CNN architecture effectively predicts downlink channel state information by learning temporal, spatial, and frequency correlations from past channel sequences. [31] proposed an ML-based architecture for estimating channels by using CNN-AR, which achieves gains in prediction quality.
@@ II-C CSI Classification
In wireless sensing, many tasks can be framed as CSI classification, including gesture recognition[32], fall detection [4], and activity recognition [33]. Reviewing CSI basics, human movement in the environment can cause changes in the multi-path channel, which in turn changes the extracted CSI matrix. Therefore, by analyzing the pattern of changes in CSI, it is possible to sense human activities.
In real-world sensing systems, accurately characterizing human actions proves to be challenging. Complex actions, such as falling, kicking, and squatting, are particularly difficult to model mathematically. Additionally, noise in Wi-Fi extraction devices, including carrier frequency offset and sampling time offset, is hard to completely eliminate. These two challenges motivate the use of deep learning techniques for Wi-Fi sensing. In [34], the authors estimate human poses from videos to supervise the training of Wi-Fi CSI. The broad learning system proposed in [35] aims to achieve fast and accurate Wi-Fi fingerprint localization. Furthermore, [36] introduces Wi-Mesh, a system that utilizes Wi-Fi signals and 2D angle-of-arrival (AoA) estimation to create 3D human meshes using deep learning, demonstrating effectiveness across various environments and conditions.
However, due to the instability of actual network environments, the receiver may struggle to successfully demodulate received signals due to several factors, including weak signal strength, frequency interference, and hardware errors. Moreover, packet loss itself may contain valuable sensing information. For instance, when an object obstructs the line-of-sight (LOS) path, the signal strength at the receiver (RX) significantly decreases, leading to an increase in packet loss. The discontinuous CSI resulting from packet loss can severely degrade the performance of wireless sensing models. Unfortunately, this issue has not been widely addressed in existing literature. Most studies attempt to use linear interpolation when the packet loss rate is low [37]. However, when the packet loss rate significantly increases, traditional interpolation fails to accurately reconstruct the true CSI, resulting in biased parameter estimation and incorrect inference.
To tackle this challenge, our previous work introduced CSI-BERT1 [24], a self-supervised learning method designed to capture the relationships within CSI data and recover lost packets. The recovered CSI can significantly enhance the capacity of sensing models. In this work, we further optimize the structure of CSI-BERT1 to improve its ability to capture temporal information, enabling it to successfully perform CSI classification tasks using the discontinuous CSI for training. Experimental results indicate that CSI-BERT2 achieves state-of-the-art performance in CSI classification tasks, greatly surpassing both CSI-BERT1 and other existing models. Consequently, there is no need to recover CSI separately for training other models. Moreover, with the aid of our proposed time embedding layer, CSI-BERT2 can effectively handle scenarios where the testing set and training set have different sampling rates.
@@ III Methodology
Figure 3: Workflow: Similar to most pre-trained models, CSI-BERT2 also has three phases: pre-training, fine-tuning, and inference. In the figure, the [MASK] tokens with green background represent randomly added [MASK] tokens, the [MASK] tokens with blue background represent [MASK] tokens used to replace [PAD], the [MASK] tokens with red text represent the [MASK] tokens that the model is expected to recover or predict accurately, and the [MASK] tokens with black text represent [MASK] tokens whose corresponding outputs are not of concern.
@@ III-A Overview
The proposed CSI-BERT2 model is a BERT-based approach designed for CSI prediction, and classification, as illustrated in Fig. 1. The workflow of our model includes three phases: pre-training, fine-tuning, and inferring, as shown in Fig. 3. During pre-training, the CSI sequence is randomly destructed, and the model is trained to recover the missing components in a self-supervised manner, to make the model learn the CSI pattern and inner relationship. After that, the model can be fine-tuned for prediction tasks or classification tasks. Finally, the trained model can be used to infer for specific downstream tasks.
@@
In this section, we leverage the CSI sequence $c=[c_{1},c_{2},...,c_{n}]$ and the corresponding timestamps $t=[t_{1},t_{2},...,t_{n}]$ as input. Each $c_{i}$ represents a flattened vector of the CSI matrix at time $t_{i}$, which contains the amplitude and phase information of subcarriers.
III-B Data Process

### v2

# CSI-BERT2: A BERT-inspired Framework for Efficient CSI Prediction and Classification in Wireless Communication and Sensing

@@ Abstract
Channel state information (CSI) is a fundamental component in both wireless communication and sensing systems, enabling critical functions such as radio resource optimization and environmental perception. In wireless sensing, data scarcity and packet loss hinder efficient model training, while in wireless communication, high-dimensional CSI matrices and short coherent times caused by high mobility present challenges in CSI estimation. To address these issues, we propose a unified framework named CSI-BERT2 for CSI prediction and classification tasks. Building on CSI-BERT, we introduce a two-stage training method that first uses a mask language model (MLM) to enable the model to learn general feature extraction from scarce datasets in an unsupervised manner, followed by fine-tuning for specific downstream tasks. Specifically, we extend MLM into a mask prediction model (MPM), which efficiently addresses the CSI prediction task. We also introduce an adaptive re-weighting layer (ARL) to enhance subcarrier representation and a multi-layer perceptron (MLP)-based temporal embedding module to mitigate permutation invariance issues in time-series CSI data. This significantly improves the CSI classification performance of the original CSI-BERT model. Extensive experiments on both real-world collected and simulated datasets demonstrate that CSI-BERT2 achieves state-of-the-art performance across all tasks. Our results further show that CSI-BERT2 generalizes effectively across varying sampling rates and robustly handles discontinuous CSI sequences caused by packet loss—challenges that conventional methods fail to address.
@@ I Introduction
Channel state information (CSI) plays a critical role in both wireless sensing and wireless communication systems by capturing the propagation characteristics of the wireless channel. This information provides valuable insights into the radio environment, including the channel conditions, signal quality, and the motion and location of objects. In wireless communication systems, CSI is essential for optimizing key processes in wireless communication systems, such as channel compensation, adaptive modulation and coding, user selection and scheduling. Effective utilization of CSI enables improved system performance, increased spectrum efficiency, and more reliable communication, ultimately enhancing the overall user experience [1, 2, 3]. In wireless sensing applications, CSI can enhance the ability to detect and track objects, leading to improved environmental awareness. Owing to its advantages in privacy protection, low cost, and penetration ability, wireless sensing has been widely adopted in diverse applications, such as fall detection [4], localization [5], and people identification [6].
Despite the importance of CSI in wireless systems, acquiring, processing, and utilizing CSI data poses significant challenges. In wireless communication, estimating CSI is time-consuming, particularly in high-dimensional and high-mobility scenarios. Frequent updates in rapidly changing environments increase computational complexity. Consequently, delays in obtaining accurate CSI can degrade spectrum efficiency by leading to suboptimal resource allocation and increased interference. In wireless sensing, packet loss due to factors such as channel noise and device errors may severely impact model performance. As noted in [7], even in relatively simple indoor environments, the average packet loss rate can reach 14.5%, with peaks up to 70% within a single second. Such discontinuities in CSI sequences pose serious challenges for the design and deployment of wireless sensing models. Additionally, for data-driven methods in both scenarios, a major challenge is the limited availability of datasets. Collecting large datasets independently is costly, while public datasets are often hetergeneous and scarce, making them difficult to use in combination. Consequently, how to enable models to learn effectively from limited datasets has become an important research topic.
Recent studies have noted and addressed these challenges. In wireless communication, the high computational complexity of CSI estimation has driven the adoption of deep learning techniques to predict future CSI more efficiently. However, most existing approaches rely on auto-regressive (AR) manners [8, 9, 10, 11], which have been criticized for error accumulation [12]. Regarding the packet loss problem in wireless sensing, most previous methods employ interpolation techniques such as linear interpolation and Kriging to estimate CSI within each dimension, often neglecting the inter-dependencies among subcarriers. Prior research [7] has indicated that these interpolation methods provide only limited benefits for sensing models. To address the issue of data scarcity, the pre-training and fine-tuning paradigm has been extensively explored in wireless systems and has shown promising performance [13]. By pre-training on either downstream datasets or unlabeled datasets, the model can achieve a more generalized capacity for data representation, thereby enhancing its generalization and robustness in downstream tasks. [^1] To avoid confusion, we clarify the concept of “pre-training” in this paper. We use this term in its original and broadest sense: any initial training phase that improves model performance on downstream tasks, regardless of the data source (labeled/unlabeled, in-domain/out-of-domain) or methodology (self-supervised, supervised, etc.). This intentionally differs from narrower interpretations that restrict “pre-training” exclusively to large-scale unsupervised learning on massive datasets, as commonly employed in today’s large models. Finally, given the wide diversity of tasks in wireless communication and sensing, designing specialized models for each task can be labor-intensive and inefficient. A unified and generic framework is therefore highly desirable [14], and many works [15, 16, 17] have begun exploring the development of foundational models for wireless systems.
To address the aforementioned shortcomings, we propose a method inspired by bidirectional encoder representations from transformers (BERT) [18], named CSI-BERT2, which is an improved version of our previous work, CSI-BERT [7]. In CSI-BERT, the masked language model (MLM) approach was first introduced for the pre-training of the CSI recovery model in Wi-Fi sensing, enabling it to learn effectively from limited dataset and improving performance in downstream tasks. Building on this, CSI-BERT2 introduces the mask prediction model (MPM), where we mask the final portion of the CSI sequence and train the model to recover it. By predicting the masked CSI in a single step, our method efficiently avoids the error accumulation problem common in AR models. Additionally, we propose a novel multi-layer perceptron (MLP)-based time embedding method, effectively addressing the permutation invariance problem [19]. Unlinke traditional positional embeddings in Transformers that captures only distance relationships, our proposed time embedding method explicitly exploit the temporal order, which is critical for CSI sequences. This design enables CSI-BERT2 to generalize efficiently in environments with varying sampling rates, and perform robustly even in a zero-shot setting. Finally, considering the varying information content convened across different subcarriers, we introduce an adaptive re-weighting layer (ARL) [20], which helps to adaptively capture the importance of each subcarrier. Through these architectural enhancements, we successfully address the low performance of the original CSI-BERT in CSI classification tasks. In conclusion, the main contributions of this paper are summarized as follows:
@@
(1) A Multifunctional Framework for CSI Time Series: We propose CSI-BERT2, a multifunctional model for CSI time series tasks in wireless systems, including CSI prediction and CSI classification. The model supports easy adaptation to various tasks simply by replacing the classification or prediction head. To realize these, we optimize the structure and training process of the original CSI-BERT model, expanding its application from solely CSI recovery to include both CSI classification and prediction.
@@
(2) Improved Model Structure: We design a novel time embedding technique that utilizes MLP to further encode temporal information, addressing the permutation invariance problem inherent in the Transformer positional embedding layer. This design enables our model to function effectively in various sampling rate scenarios. Furthermore, we propose a simple attention mechanism called ARL to better capture the relationships and importance between subcarriers adaptively.
(3) Improved Training Process: To make full use of the limited training data, we propose a two-stage training strategy. Specifically, we first pre-train the model to learn general data representations using unsupervised MLM, followed by fine-tuning for specific downstream tasks. Additionally, we expand MLM to the MPM, where only the last portion of CSI sequences is masked, enabling the model to effectively perform CSI prediction tasks.
(4) Experiment Evaluation: To validate the performance of CSI-BERT2, we conduct comprehensive experiments across four datasets, including both public and self-collected data from real-world and simulated environments. In the CSI prediction task, CSI-BERT2 achieves the lowest error, outperforming the current state-of-the-art (SOTA) by over 60% in mean squared error (MSE). Although it requires slightly more computation time than other models, the total prediction time for the entire dataset, consisting of hundreds of samples, remains under 0.5 seconds, demonstrating its practicality for real-time deployment. In the CSI classification task, CSI-BERT2 consistently achieves the highest accuracy across various applications, including gesture recognition, people identification, fall detection, and crowd counting. Notably, our method shows strong robustness to discontinuous data caused by packet loss and to sampling rate mismatches between the training and testing sets, without the need to process and recover data or train a new model separately. Its accuracy drops by no more than 2%, in contrast to other models that suffer tens of percentage points in degradation. By demonstrating robust performance under packet loss and variable sampling rates, CSI-BERT2 eliminates the need for resource-intensive retraining in dynamic environments, making it viable for real-time applications such as emergency fall detection.
@@
The structure of this paper is organized as follows. Section II introduces the fundamental principles of CSI, the background on CSI prediction and classification, and a review of related works. Section III provides a detailed description of the model structure and the training process for different tasks. Section IV presents a series of comparative experiments to demonstrate the advantages of our method, and several ablation studies to illustrate the efficiency of our design. Finally, Section V summarizes the paper and highlights potential directions for future research.
II Preliminaries and Related Works
@@ II-A CSI Basics
Following [21], in Wi-Fi communication, CSI characterizes the way Wi-Fi signals propagate from the transmitter (TX) to the receiver (RX). Modern Wi-Fi systems typically employ orthogonal frequency division multiplexing (OFDM) as the modulation technique. OFDM divides the wideband channel into multiple narrowband subcarriers in the frequency domain, each experiencing flat fading for easier processing. In the time domain, data is transmitted as a series of OFDM symbols with a cyclic prefix to prevent interference from multipath delays. The amplitude and phase of CSI are influenced by multipath effects, which include amplitude attenuation and phase shift. Each element of the CSI matrix $H$ is given by:
@@ II-B CSI Prediction
Obtaining accurate CSI matrix in wireless communication remains a challenging task [22]. The large number of antennas results in a high-dimensional CSI matrix, which in turn makes the channel estimation algorithms difficult to solve. Additionally, when objects in the environment move at high speeds, the channel coherence time shortens considerably, imposing stricter requirements on the channel estimation speed. As illustrated in Fig. 2b, CSI plays a critical role in optimizing wireless communication system performance, making fast and reliable CSI estimation essential. Furthermore, predicting future CSI can provide valuable opportunities for system-level optimizations, such as effective channel compensation.
Recently, some works have explored the use of additional information for CSI prediction. For example, in [23], the authors used multimodal sensory data, including received pilots, user position, and historical downlink CSI to predict the current CSI. In [24], a deep learning-based framework using a three dimensional (3D) CNN architecture effectively predicts downlink CSI by learning temporal, spatial, and frequency correlations from past channel sequences. The authors in [25] have proposed an machine learning (ML)-based architecture for estimating channels by using CNN-AR, which achieves gains in prediction quality.
@@ II-C CSI Classification
In wireless sensing, many tasks can be formulated as CSI classification problems, including gesture recognition[26], fall detection [4], and activity recognition [27]. Reviewing CSI basics, human movement in the environment can cause changes in the multipath channel, which in turn changes the extracted CSI matrix. Therefore, by analyzing the pattern of changes in CSI, it is possible to sense human activities.
However, in real-world sensing systems, accurately characterizing human actions are challenging. Complex actions, such as falling, kicking, and squatting, are particularly difficult to model mathematically. Additionally, noise in Wi-Fi extraction devices, including carrier frequency offset and sampling time offset, is hard to completely eliminate. These challenges motivate the use of deep learning techniques for Wi-Fi sensing. For example, in [28], the authors have estimated human poses from videos to supervise the training of Wi-Fi CSI. The broad learning system proposed in [29] aims to achieve fast and accurate Wi-Fi fingerprint localization. Furthermore, [30] have proposed Wi-Mesh, which utilizes Wi-Fi signals and 2D angle-of-arrival (AoA) estimation to create 3D human meshes using deep learning, showing robustness across various environments and conditions.
Nonetheless, unstable network conditions often lead to signal demodulation failures at the receiver due to factors, such as weak signal strength, frequency interference, and hardware errors. Moreover, packet loss itself may contain valuable sensing information. For instance, when an object obstructs the line-of-sight (LoS) path, the signal strength at the receiver may significantly decreases, leading to an increase in packet loss. These discontinuities in CSI resulting from packet loss can severely degrade the performance of wireless sensing models. Unfortunately, this issue has not been well addressed in existing literature. Most studies attempt to use linear interpolation when the packet loss rate is low [31]. However, traditional interpolation methods fail under high loss rates, resulting in biased parameter estimation and incorrect inference.
To tackle this challenge, our previous work have introduced CSI-BERT [7], a self-supervised learning method designed to capture the internal structure of CSI data and recover lost packets. The recovered CSI was shown to substantially enhance the performance of downstream sensing tasks. In this work, we further optimize the structure of CSI-BERT to improve its ability to capture temporal information, enabling it to successfully perform CSI classification tasks using the discontinuous CSI for training.
@@ III-A Overview
In this section, we will introduce the data preprocessing methods, neural network architecture, pre-training process, and the fine-tuning and inference methods of CSI-BERT2 in turn. As a continuation of CSI-BERT, we first provide a brief introduction to it. (This section will primarily focus on the modified and enhanced components, while more details about CSI-BERT can be found in our previous paper [7].)
Prior to CSI-BERT, several works had already applied BERT [18] to wireless sensing tasks such as people localization [32, 33] and radio map construction [34], based on the intuition that the bidirectional encoder of the Transformer [35] can effectively capture the relationships among packets by the self-attention mechanism. However, these studies typically fed discrete signals into the original BERT embedding layer, which leads to significant information loss. In contrast, CSI-BERT first proposed a novel embedding layer tailored to the structure of CSI data, where a multi-layer perceptron (MLP) is used to encode continuous CSI data directly. Additionally, a positional-embedding style time-embedding layer was introduced to capture the temporal information among packets. Inspired by the similarity between packet recovery and masked language modeling (MLM), CSI-BERT successfully achieves high-quality packet recovery through unsupervised learning without labeled data. The recovered CSI data enhances the performance of other sensing models in classification tasks, outperforming traditional interpolation methods that fail to consider the internal relationships among subcarriers.
However, the original CSI-BERT does not fully leverage the advantages of MLM, resulting in subpar performance in downstream tasks. For example, the classification accuracy of CSI-BERT after fine-tuning does not surpass that of the vanilla ResNet18 [36]. This limitation can be attributed to the inadequacies of the embedding layer. To address this issue, CSI-BERT2 introduces an ARL-based spatial encoding module to capture the relationships and importance among subcarriers. Additionally, a MLP-based temporal encoding module is proposed to capture both the order and distance relationships among packets, effectively solving the permutation invariance problem inherent in conventional positional embedding modules, which capture distance relationships without considering order information. Moreover, to fully utilize the learned knowledge during pre-training, we propose a novel fine-tuning method named mask prediction model (MPM) for the CSI prediction task, wherein the last portion of the CSI sequence is randomly masked, and the model is trained to predict these masked values. Compared to conventional AR-based methods, our approach can complete the prediction in one step, thereby eliminating the error accumulation problem.
@@
The model structure and working scenarios of CSI-BERT2 are illustrated in Fig. 1 and Fig. 2, respectively. We leverage the CSI sequence $c=[c_{1},c_{2},\ldots,c_{N}]$ and the corresponding timestamps $t=[t_{1},t_{2},\ldots,t_{N}]$ as input. Each $c_{i}$ represents a flattened vector of the CSI matrix at time $t_{i}$, which contains the amplitude and phase information of the subcarriers.
Figure 1: CSI-BERT2 network architecture: In the figure, ‘T’ represents token, ‘t’ denotes timestamp, and ‘P’ indicates the position. The red star highlights the different components of the CSI-BERT.
(a) CSI Recovery
(b) CSI Prediction
(c) CSI Classification
Figure 2: Three tasks in this paper: (a) During the unsupervised training phase, we randomly mask some CSI tokens and train the CSI-BERT2 to recover them in an unsupervised manner, while a discriminator is employed to enhance the realism of the recovered results. (b) The CSI prediction task focuses on swiftly predicting future CSI series for proactive communication optimization. (c) The CSI classification task utilizes CSI for specific sensing applications, such as gesture recognition. Notably, our CSI-BERT2 can efficiently process CSI data across varying sampling rates, a shortcoming for most other models.
III-B Data Pre-processing